### v2

# Lightweight posterior construction for gravitational-wave catalogs with the Kolmogorov-Arnold network

Abstract
Neural density estimation has seen widespread applications in the gravitational-wave (GW) data analysis, which enables real-time parameter estimation for compact binary coalescences and enhances rapid inference for subsequent analysis such as population inference. In this work, we explore the application of using the Kolmogorov-Arnold network (KAN) to construct efficient and interpretable neural density estimators for lightweight posterior construction of GW catalogs. By replacing conventional activation functions with learnable splines, KAN achieves superior interpretability, higher accuracy, and greater parameter efficiency on related scientific tasks. Leveraging this feature, we propose a KAN-based neural density estimator, which ingests megabyte-scale GW posterior samples and compresses them into model weights of tens of kilobytes. Subsequently, analytic expressions requiring only several kilobytes can be further distilled from these neural network weights with minimal accuracy trade-off. In practice, GW posterior samples with fidelity can be regenerated rapidly using the model weights or analytic expressions for subsequent analysis. Our lightweight posterior construction strategy is expected to facilitate user-level data storage and transmission, paving a path for efficient analysis of numerous GW events in the next-generation GW detectors.
I Introduction
Since the first detection of the gravitational-wave (GW) signal from a binary black hole (BBH) merger LIGOScientific:2016aoc, GWs have opened a new era in our exploration of the Universe. The ground-based GW detectors, LIGO LIGOScientific:2014pky, Virgo VIRGO:2014yos, and KAGRA KAGRA:2018plz, have completed three observing runs LIGOScientific:2018mvr; LIGOScientific:2020ibl; LIGOScientific:2021usb; KAGRA:2021vkt, and together with the first part of the fourth run LIGOScientific:2025hdt; LIGOScientific:2025slb, more than 200 GW events from compact binary coalescences (CBCs) are reported. These observations have driven transformative advances across multiple domains of astrophysics LIGOScientific:2018jsj; LIGOScientific:2020kqk; KAGRA:2021duu, cosmology ligo2017gravitational; DES:2019ccw; LIGOScientific:2019zcs; LIGOScientific:2021aug, and fundamental physics Yagi:2013awa; LIGOScientific:2018cki; LIGOScientific:2021sio.
Behind these groundbreaking discoveries lie the advancements in instruments, as well as data analysis techniques. Bayesian inference constitutes the fundamental framework in GW data analysis, providing a principled and systematic approach to parameter estimation and population inference thrane2019introduction. In Bayesian inference, the parameter estimation is calculating the conditional distribution, the so-called posterior $P(\bm{\theta}|d)$, of the parameters $\bm{\theta}$ given the observed data $d$. By employing Bayes’ theorem, the posterior can be obtained given the prior $\pi(\bm{\theta})$ and the likelihood function $\mathcal{L}(d|\bm{\theta})$, via $P(\bm{\theta}|d)={\mathcal{L}(d|\bm{\theta})\pi(\bm{\theta})}/{\mathcal{Z}(d)}$, where $\mathcal{Z}(d)$ is the normalization factor known as the evidence. Usually, it is unfeasible to directly calculate the exact expression of the posterior. Luckily, with advanced stochastic sampling methods such as Markov Chain Monte Carlo (MCMC) Christensen:1998gf; Sharma:2017wfu and nested sampling Skilling:2004pqw; Skilling:2006gxv, we can draw posterior samples from this distribution. Posterior samples constitute key data products in GW catalogs underpinning subsequent scientific investigation in GW astronomy.
As GW detections become routine, the focus of GW astrophysics is expanding from analyzing individual events to conducting systematic studies on the entire population of sources. This shift necessitates a compilation of comprehensive and accessible GW catalogs. Growing catalogs call for powerful statistical techniques, like hierarchical Bayesian analyses Mandel:2018mve; thrane2019introduction; Vitale:2020aaz; Wang:2024xon; Gerosa:2024isl, to unravel the collective properties of compact binaries. Distributions of compact binary source parameters, such as masses KAGRA:2021duu; Golomb:2021tll; Dong:2024zzl and spins KAGRA:2021duu; Zhu:2017znf; Tong:2022iws provide critical constraints on models of stellar evolution and compact object formation. Furthermore, combining information from multiple events allows for stringent tests of Einstein’s theory of general relativity LIGOScientific:2021sio and precise constraints on the properties of dense matter, such as the equation of state of neutron stars Landry:2018prl; Wysocki:2020myz; Shao:2022koz; Wang:2024xon. Therefore, efficiently constructing and utilizing GW catalogs is crucial to advance our understanding of the Universe, especially in the era of next-generation GW detectors when many more events are detected.
Next-generation ground-based GW detectors, such as the Einstein Telescope (ET) Punturo:2010zz; Branchesi:2023mws; Abac:2025saz and Cosmic Explorer (CE) Reitze:2019iox; Evans:2021gyd, will achieve unprecedented sensitivities, unlocking new scientific opportunities Sathyaprakash:2019yqt; Kalogera:2021bya. They will boost the detection rates for binary neutron-star and black-hole mergers, reaching $10^{5}$$10^{6}$ events per year Himemoto:2021ukb; Samajdar:2021egv; Hu:2022bji; Borhanian:2022czq; Pieroni:2022bbh; Ronchini:2022gwk; Iacovelli:2022bbs; Begnoni:2025oyd. Considering that standard offline Bayesian inference for a single event already demands high computational resources, the high event rates expected in the next-generation GW detectors will introduce new computational challenge for data analysis. As predicted, performing standard Bayesian inference on a one-month catalog requires $10^{13}$$10^{15}$ CPU hours, imposing a substantial burden on computing infrastructure, electricity cost, and environmental impact Hu:2024mvn. Efficient, accelerated Bayesian methods are therefore urgently needed. Besides, from the perspective of data storage and transmission, posterior-sample-based GW catalogs with millions of events could easily reach a terabyte-scale volume, placing significant burdens on users for both transmission and storage.
The development of machine learning techniques offers a possible solution to the data challenges posed by next-generation GW detectors. Machine learning has already demonstrated its great potential in GW detection George:2017pmj; Huerta:2017kez; Xia:2020vem; Ma:2023ctz; Ma:2024log; Wang:2025xvj, parameter estimation Gabbard:2019rde; Green:2020hst; DEmilio:2021laf; Santoliquido:2025lot; Green:2020dnx; Dax:2021tsq; Dax:2024mcn, and subsequent astrophysical researches Leyde:2023iof; mcewen2021machine; Talbot:2020oeu; Mould:2022ccw; Ventagli:2024xsh; Tang:2024orq; Hu:2025vlp. Specifically, neural posterior estimation has been widely adopted across multiple GW Bayesian inference tasks Dax:2021tsq; Dax:2024mcn; Leyde:2023iof; mcewen2021machine, accelerating the analysis by orders of magnitude. In particular, neural posterior estimation with normalizing flows Green:2020dnx; Dax:2021tsq; Dax:2024mcn; Leyde:2023iof; Langendorff:2022fzq; kobyzev2020normalizing achieves real-time Bayesian inference for binary black hole mergers and binary neutron star mergers, opening new avenues for multi-messenger astronomy. In the era of next-generation detectors, neural posterior estimation has also been explored to address data challenges such as overlapping signals Langendorff:2022fzq; Hu:2025vlp and long-duration signal analysis Hu:2024lrj. For neural posterior estimation, it needs a neural-network-based density estimator to represent distributions. These neural density estimators are usually built on generative-model architectures, such as variational autoencoders Kingma:2013hel; Gabbard:2019rde, normalizing flows, and flow matching frameworks Chen:2018wjc; lipman2023flow; Liang:2024new. For instance, normalizing flows learn a sequence of invertible transformations that map a simple base distribution to the complex target distribution, enabling both fast sampling and exact density evaluation. By efficiently capturing high-dimensional, multimodal distributions, neural density estimators show great promise for alleviating the computational bottlenecks in GW Bayesian inference.
In this work, we propose a Kolmogorov-Arnold network (KAN) liu2025kan; liu2024kan2 based neural density estimator and explore its potential for constructing high-fidelity, low-storage surrogate GW catalogs. KAN is a machine-learning architecture inspired by the Kolmogorov-Arnold representation theorem arnold2009representation; braun2009constructive. Rather than using fixed activations in conventional multi-layer perceptrons (MLPs), KAN employs edge-wise learnable activation functions. This modification enables KAN to achieve superior interpretability, higher accuracy, and greater parameter efficiency compared to conventional architectures on small-scale artificial intelligence (AI) for scientific tasks. Furthermore, this design also allows the network to converge toward interpretable analytic expressions, a process referred to as the symbolification of the neural network. Currently, KANs are used in fields such as fluid dynamics Toscano2025; Wang_2025, nuclear physics PhysRevC.111.024316, complex dynamical systems PhysRevResearch.7.023037, molecular property prediction li2025KAGNN, and cosmology Cui:2025rri; Jones:2025obf, and KAN-based machine learning architectures are also rapidly advancing yang2025kolmogorovarnold; fang2025kaa; huang2025timekan. The rapid development of KAN offers a promising approach to enhance the interpretability of neural networks in AI for science tasks. In our case, to support downstream studies based on GW posteriors, one can download compact, encoded files from our surrogate GW catalogs and rapidly resample the posterior distributions locally. This strategy enables efficient catalog construction while dramatically reducing the demands on data transmission and storage.
Leveraging KAN’s efficient representation and unique symbolification capability, our KAN-based neural density estimator can provide two compressed representations of GW posterior samples. The first is the set of neural network weights, which prioritize high fidelity to support precision-critical analysis tasks. The second representation, analytic fits of the posterior distributions, maximizes compression by encoding the posteriors in compact, closed-form analytic expressions, though with a small trade-off in accuracy. Both products enable fast posterior resampling, and their reliability is validated in downstream analysis tasks through GW population inference, exhibiting great potential in efficient and robust GW catalog construction for future GW observations.
The paper is organized as follows. In Sec. II, we introduce the machine learning techniques underlying the KAN-based neural density estimator and present the flowchart for efficient GW catalog construction. We showcase three verifiable simple examples in Sec. III to demonstrate the feasibility of the proposed method, validating the effectiveness and accuracy of analytic probability density function (PDF) fitting using the KAN network. In Sec. IV, we present the compression results based on the KAN-based neural density estimator on posterior samples from some current GW events and further validate the reliability of the compact data products by performing population inference tasks using the resampled samples from the compressed data products. Finally, Sec. V summarizes our findings and discusses potential directions for future improvements.
Figure 1: Flowchart for lightweight GW catalog construction. Starting from raw posterior samples of a single GW event as our training set, we first train the KAN-based neural density estimator to minimize the average negative log-posterior. As shown in the box at the bottom left, we replace MADE’s unmasked edges with KAN’s learnable nonlinear edges (a linear combination of B-spline curves). Once the loss has converged, the analytic expression of each unmasked edge is obtained and synthesized through the network topology to get analytic expressions of all output nodes: $\ln(a_{ij})$, $\mu_{ij}$, and $2\ln(\sigma_{ij})$; see main text for details. Expressions of all output nodes are combined to obtain the analytic joint PDF. After training and symbolification, we store the neural network weights, typically of size $\mathcal{O}\,(10^{1}\,{\rm KB})$, and analytic expressions of output nodes, typically of size $\mathcal{O}\,(1\,{\rm KB})$, as compact data products. Users can download these data products and use them to compute probability density or resample posterior samples for downstream analysis tasks.
II KAN-based neural density estimator
In this section, we introduce the method to compress the posterior samples through a KAN-based neural density estimator. First, we introduce KAN liu2025kan; liu2024kan2 in Sec. II.1. We explain how KAN improves upon conventional neural networks and enables symbolification, after which the entire model is represented merely by the analytic expressions of its output nodes. In Sec. II.2, we present the Masked Autoencoder for Distribution Estimation (MADE) germain2015made, one of the core building blocks commonly employed in normalizing flows Papamakarios:2017tec. MADE serves as our base density estimator architecture, which we then enhance using the KAN framework. Section II.3 describes how to perform rapid resampling using the neural network weights and/or the analytic expressions. We outline the workflow for lightweight GW catalog construction in Fig. 1.
II.1 Kolmogorov-Arnold Network
KAN is a novel machine-learning architecture inspired by the Kolmogorov-Arnold representation theorem arnold2009representation; braun2009constructive. Unlike conventional MLPs that rely on fixed activation functions at each neuron, KAN introduces learnable activation functions on the edges connecting neurons between layers. These edge-wise functions are parameterized as a linear combination of B-spline curves gordon1974b, with coefficients being learnable parameters that are optimized during training. This design allows KAN to capture complex nonlinear relationships between inputs and outputs more flexibly, leading to enhanced accuracy and greater parameter efficiency, particularly in small-scale scientific AI applications.
A key feature of KAN is its capability for network symbolification, a two-stage process involving function matching and synthesis. First, the trained B-spline activation functions on each edge are matched to a predefined library of elementary mathematical functions (e.g., polynomial, trigonometric, and exponential functions) to derive their analytic representations. Subsequently, these individual analytic expressions are composed according to the network’s topology, yielding closed-form analytic expressions that describe the network’s outputs as explicit functions of its inputs liu2025kan. This process significantly enhances the interpretability of the neural network, making its internal mechanism more transparent and analyzable. Leveraging KAN’s efficient function representation and unique symbolification capabilities, we propose a KAN-based neural density estimator. Our objective is to harness these advantages to achieve substantial storage compression of GW posterior samples, thereby offering a new approach for constructing lightweight and user-friendly GW catalogs.
II.2 Masked Autoencoder for Distribution Estimation
To model the probability distribution of the posterior samples, we employ the MADE framework. MADE facilitates density estimation by decomposing the joint probability distribution autoregressively germain2015made. For a $D$-dimensional parameter vector $\bm{\theta}$, this decomposition is expressed as
$P(\bm{\theta})=\prod_{i=1}^{D}P_{i}(\theta_{i}|\bm{\theta}_{<i})\,$, (1)
where $\bm{\theta}_{<i}$ denotes the set of parameters $\{\theta_{1},\theta_{2},\cdots,\theta_{i-1}\}$. MADE approximates each conditional probability $P_{i}$ with a Gaussian mixture model (GMM) mcewen2021machine; germain2015made; papa2016fast, such that
$P_{i}(\theta_{i}|\bm{\theta}_{<i})\approx\sum_{j=1}^{G}a_{ij}(\bm{\theta}_{<i})\mathcal{N}\left(\theta_{i}\,\big|\,\mu_{ij}(\bm{\theta}_{<i}),\sigma_{ij}^{2}(\bm{\theta}_{<i})\right)\,.$ (2)
Here, $a_{ij}$, $\mu_{ij}$, and $\sigma_{ij}$ are the mixture weights, means, and standard deviations of the Gaussian components, respectively, all of which are functions of the preceding parameters $\bm{\theta}_{<i}$. The number of Gaussian components used in each GMM is denoted as $G$.
A neural network, denoted by its weights $w$, is trained to output the parameters of these GMMs. To ensure positivity of $a_{ij}$ and $\sigma_{ij}$, the network predicts $\ln(a_{ij})$ and $2\ln(\sigma_{ij})$ alongside $\mu_{ij}$. The complete network-approximated distribution is denoted as $P_{w}(\bm{\theta})$. The crucial autoregressive property is enforced by applying binary masks to the network’s connections, ensuring that each conditional probability $P_{i}$ depends only on $\bm{\theta}_{<i}$. Alternative approaches for enforcing autoregressive constraints include masked convolutions van2016pixel and causal convolutions oord2016wavenet. If the target distribution is conditional $P(\bm{\theta}|\bm{\phi})$, the conditioning variables $\bm{\phi}$ need to be connected to the network through unmasked connections.
In our implementation, we enhance the standard MADE architecture by replacing its linear layers with KAN layers, creating a KAN-based MADE. This allows us to later extract analytic expressions for the PDF. Leveraging KAN’s parameter efficiency, our framework is expressive enough even with a light network structure and also only requires small storage space. The network is trained by minimizing the average negative log-posterior, $\mathbb{E}_{P(\bm{\theta})}\big[\ln P_{w}(\bm{\theta})\big]$, using raw posterior samples as the training set. After loss convergence, we obtain two compact data products:
  1. (i)
    Neural Network Weights: The trained weights $w$ of the KAN-based MADE, which provide a high-fidelity representation of the posterior distribution.
  2. (ii)
    Analytic Expressions: By applying KAN’s symbolification process, we derive analytic expressions for all GMM parameters ($a_{ij}$, $\mu_{ij}$, $\sigma_{ij}$) as functions of $\bm{\theta}_{<i}$. These expressions offer a maximally compressed, interpretable representation of the posterior. We also offer functionality to express the joint probability distribution $P_{w}(\bm{\theta})$ in closed form using these saved expressions according to Eq. (1) and Eq. (2), in order to facilitate subsequent probability density evaluation.
These data products enable users to reconstruct the posterior distribution for further analysis with little storage and computational overhead, compared with the original posterior samples in the GW catalogs.
II.3 Resampling Using Neural Network Weights and Analytic Expressions
Both the neural network weights and the analytic expressions can be used to generate new posterior samples and evaluate the probability density. The resampling process is autoregressive: for each dimension $i=1,\dots,D$, $\theta_{i}$ is sampled from the conditional distribution $P_{i}(\theta_{i}|\bm{\theta}_{<i})$ using the already-sampled values of $\bm{\theta}_{<i}$. The resampling process using analytic expressions is more efficient than resampling with the neural network weights.
For the resampling with neural network weights, the sample generation process is accomplished iteratively according to Eq. (1) where $\bm{\theta}$ is set to $\bm{0}$ in the beginning. During the $i$-th iteration, $\bm{\theta}_{<i}$ is already generated and $\bm{\theta}_{\geq i}$ is still $\bm{0}$. Then $\{\bm{\theta}_{<i},\bm{0}\}$ is fed into the network to obtain $a_{i,1\leq j\leq G}$, $\mu_{i,1\leq j\leq G}$, and $\sigma_{i,1\leq j\leq G}$. Next, $\theta_{i}$ is sampled from the conditional distribution $P_{i}(\theta_{i}|\bm{\theta}_{<i})$. However, in each sequential iteration, all network outputs are computed, introducing computational inefficiency as many node computations—those not related to $P_{i}$—are unnecessary. If we assume that the network does not contain hidden layers (as we do), then the computational complexity of resampling is $\mathcal{O}\,\big(GD^{3}\big)$. This is because if the number of nodes in the input layer is $D$, the number of nodes in the output layer is proportional to $GD$, and the number of iterations equals $D$. Probability density computation using a neural network requires one forward computation to get all the GMM parameters and the complexity is $\mathcal{O}\,\big(GD^{2}\big)$. Then $P_{w}(\bm{\theta})$ is computed according to Eq. (1) and Eq. (2).
For the resampling with analytic expressions, it enables more efficient sampling through direct parameter computation. Specifically, in each iteration, the first i1i-1 dimensions can be directly substituted into analytic expressions of $a_{i,1\leq j\leq G}$, $\mu_{i,1\leq j\leq G}$, and $\sigma_{i,1\leq j\leq G}$. This optimized process eliminates computation of output nodes that are not related to $P_{i}$. It reduces sampling complexity to $\mathcal{O}\,(GD)$ and provides substantial efficiency gains in high-dimensional cases. Probability density computation using analytic expressions simply requires substituting observed values into the analytic expression of $P_{w}(\bm{\theta})$. This direct evaluation process requires only basic math operations. Furthermore, the analytic $P_{w}(\bm{\theta})$ provides mathematical tractability, permitting direct manipulation of the PDF through algebraic operations.
Figure 2: Results of three verifiable cases. Marginalized one- and two-dimensional distributions are shown in corner plots, comparing samples resampled from fitted expressions (orange) and samples generated from ground-truth distributions (blue). Both of them contain 40,000 samples. In Case 2, $\phi$ is fixed at 0 during the sampling process. Contour lines in the two-dimensional joint distributions delineate the $50\%$ and $90\%$ credible regions. The second row displays the ground-truth PDFs of the three cases, where $\mathcal{N}(\mu,\sigma^{2})$ denotes a Gaussian distribution with mean value $\mu$ and variance $\sigma^{2}$. The third row shows the analytic expressions of the fitted PDFs chosen by our KAN-based neural density estimator where the coefficients are rounded to two decimal places.
III Case Tests
Before applying the KAN-based neural density estimator to real GW posterior data, we construct three illustrative cases to test the reliability of our model and explore its ability to fit posterior distributions into analytic expressions. We choose a unimodal distribution, a conditional unimodal distribution, and a bimodal distribution as our test cases Papamakarios:2017tec. In each example, we start with the ground-truth PDF to obtain the raw samples, and then use these as input for the KAN-based neural density estimator to fit for the distribution’s analytic form. It is important to note that we do not specify the analytic form of the distribution in advance. Instead, we allow the KAN-based neural density estimator to freely explore the fitting expressions without any additional input information from human.
The PDFs of the three selected examples are as follows. The first example of a unimodal distribution has a PDF,
$p(\theta_{1},\theta_{2})=\mathcal{N}(\theta_{1}|0,2^{2})\times\mathcal{N}\Big(\theta_{2}\,\big|\,\frac{1}{4}\theta_{1}^{2},1\Big)\,.$ (3)
Its distribution forms a crescent shape, and we aim to use this example to test the method’s capability of accommodating non-Gaussian distributions. The PDF of the second example, a conditional unimodal distribution, reads
$p(\theta_{1},\theta_{2}\,|\,\phi)=\mathcal{N}\big(\theta_{1}|\sin\phi,0.1^{2}\big)\times\mathcal{N}\big(\theta_{2}|\cos\phi,0.1^{2}\big)\,.$ (4)
This distribution is conditional on $\phi$. Although an unconditional neural density estimator is sufficient for this work, exploring conditional density estimators is more important for future applications in the development of neural posterior estimation algorithms. The PDF of the third example, a bimodal distribution, is
$p(\theta_{1},\theta_{2})=$ $\left[\frac{1}{2}\mathcal{N}\big(\theta_{1}|-6,2^{2}\big)+\frac{1}{2}\mathcal{N}\big(\theta_{1}|6,2^{2}\big)\right]$ (5)
$\times\left[\frac{1}{2}\mathcal{N}\Big(\theta_{2}\,\big|\,\frac{1}{4}\theta_{1}^{2},1\Big)+\frac{1}{2}\mathcal{N}\big(\theta_{2}|0,1\big)\right]\,.$
This example involves a highly distorted bimodal distribution, which is used to test the neural density estimator’s fitting ability with multimodal distributions. Beyond these examples, we also present one additional illustrative example based on an exponential distribution in Appendix A to more comprehensively demonstrate the effectiveness of this method in modeling diverse distributions.
The verification process is nearly identical to the workflow shown in Fig. 1. We start with raw samples and obtain the encoded data products. The difference lies in the fact that in the verifiable examples, we know the ground-truth PDFs, which allows us to compare them with the analytic fitting expressions from KAN. In the tested examples, for the KAN-based MADE network architecture, we include only the input and output layers, without any intermediate hidden layers. In all three examples, each unmasked edge of the network is set to contain eight cubic B-spline functions. We find that, due to the powerful fitting ability of the learnable spline edges in KAN, this simplest MADE design is already remarkable to support high-precision distribution fitting.
The results are shown in Fig. 2, where we present the distribution expressions fitted by the KAN-based neural density estimator in the last row, along with the graphic distribution results obtained by resampling from these fitted expressions. As we can see, even without any additional information, KAN successfully captures accurately the underlying analytic fitting form. In the applications to GW catalogs, compared to the raw samples, the encoded analytic data products offer significant advantages in terms of storage and allow for more flexibility in determining the number of resampled points to meet varying computational needs. In the analytic fitting results, the algorithm uses one Gaussian component for fitting in the first and second examples, while it uses two Gaussian components for fitting in the third example.
To further demonstrate the storage efficiency, we conduct a controlled comparison between storing samples and storing the encoded products using Case 3. Encoded analytic data products exhibit exceptional storage efficiency, requiring only $\sim 1\,{\rm KB}$ of storage. The same storage corresponds to approximately 110 samples. For comparison with other methods, we present the distribution results from expanding these 110 samples using kernel density estimation (KDE) in Fig. 3. Clearly, the posterior distribution represented by these 110 samples exhibits significant fluctuations and spurious structures, far from accurately capturing the true distribution. Our KAN-based density estimator (see the third corner plot in Fig. 2), in contrast, offers clear advantages in interpretability and storage efficiency.
Figure 3: Results of KDE for Case 3. The 110 samples are utilized in a Gaussian kernel with a bandwidth of 0.35. The KDE is accomplished using the scikit-learn package. Marginalized one- and two-dimensional distributions are shown in the corner plot, comparing samples resampled using KDE (red) and samples generated from the ground-true distribution (blue). Both of them contain 40,000 samples. Contour lines in the two-dimensional joint distributions delineate the $50\%$ and $90\%$ credible regions.
IV Lightweight posterior construction for GW catalogs
In this section, we apply the KAN-based neural density estimator to real GW data. We select six GW parameters—chirp ${\cal M}, mass$ ratio $q$, two spin magnitudes $\chi_{1}$ and $\chi_{2}$, and two spin inclinations $\theta_{1}$ and $\theta_{2}$—and attempt to encode the posterior distributions of ten GW events: GW150914, GW151012, GW151226, GW170104, GW170608, GW170729, GW170809, GW170814, GW170818, and GW170823 LIGOScientific:2021usb. Brief information of these events is presented in Appendix B. The posterior samples are obtained from the Gravitational Wave Open Science Center (GWOSC) [^1] https://gwosc.org/ LIGOScientific:2019lzm; KAGRA:2023pio; LIGOScientific:2025snk. We evaluate the reliability of our two encoded data products—high-fidelity neural network weights and analytic expressions—by comparing the posteriors of raw samples with those of the resampled samples. Additionally, we further ensure the reliability of encoded data products in practical applications by performing population inference using reconstructed posterior samples.
For the chosen GW posteriors, a key challenge for us is that some parameter distributions are non-smooth, being truncated at their prior boundaries; for example, ${\cal M}>0$ and $q\leq 1$. Direct density estimation on such truncated distributions would require substantially more network parameters to capture these sharp edges. To address this, we apply a data-driven preprocessing step that reparameterizes each truncated distribution into an untruncated form defined over the full parameter space, easing the estimation task and potentially enhancing the model’s robustness. Specifically, we remove the bounds on the parameters using invertible transformations. Because the chirp mass is limited to be positive, we apply a transform through
$\ln{\mathcal{M}}^{\prime}=\frac{\ln{\mathcal{M}}-\mu[\ln{\mathcal{M}}]}{\sigma[\ln{\mathcal{M}}]}\,$, (6)
where $\mu[\ln{\mathcal{M}}]$ and $\sigma[\ln{\mathcal{M}}]$ are the mean and standard deviation of the logarithm of the chirp mass respectively. Among the remaining parameters, the mass ratio and spin magnitudes are constrained within the interval $[0,1]$, while the spin inclinations are restricted to $0,\pi]$. We follow Hoy:2020vys and apply the logit transformation to these parameters,
$\theta^{\prime}=\ln\left(\frac{\theta-a}{b-\theta}\right)\,.$ (7)
Here, $\theta$ represents the original parameter value, and $a$ and $b$ are the lower and upper bounds of the parameter, respectively. All these transformations are invertible and performed before training the neural network. After the training phase, inverse transformations are performed during the sampling phase to recover the original parameter values. Other methods such as boundary reflection can also be used to handle the bounded parameters jones1993simple.
After data preprocessing, neural network training, and symbolification, we obtain two encoded representations of the posterior samples for each event. For these ten GW events, all neural networks adopt a compact two-layer architecture. We empirically explored different choices for the number of Gaussian components $G$ and B-spline basis functions $B$. Increasing $G$ and $B$ beyond moderate values yielded only marginal gains, whereas reducing them degraded reconstruction fidelity. Accordingly, we selected $G=5$ and $B=23$ as a practical compromise between accuracy and compactness. Taking GW150914 as an example, Tab. 1 lists the storage requirements and sampling rates for both types of encoded data products. The neural network weights and analytic expressions achieve approximately $90\times$ and $3500\times$ compression ratios in storage, respectively.
The resampling results are shown in Fig. 4 and Fig. 5. As shown in Fig. 4, the samples generated from the neural network weights exhibit close agreement with the original distribution. In Fig. 5, samples generated with the analytic expressions capture the overall structure of the posterior distribution, but exhibit some deviations from the original distributions. Compared with the neural network weights, this reflects some trade-off between fidelity and the gains in storage efficiency, interpretability, and expression simplicity. By replacing the network’s nonlinear edge functions with more restrictive analytical forms, the model’s complexity is inevitably reduced and some fidelity in capturing local features is lost. The analytic expression data products occupy only $\mathcal{O}\,(1\,{\rm KB})$ of storage. In Fig. 5 we also show samples from a fitted multivariate Gaussian distribution, where we use the same data preprocessing method as in the KAN-based neural density estimator. In preparing them, we start from the initial mean and covariance matrix of the preprocessed raw samples, and obtain the optimal mean and covariance matrix of the 6-dimensional Gaussian distribution by minimizing the average negative log-posterior. Compared to other representations, such as summarizing GW posteriors by their mean and covariance matrix, as well as its higher-order extension Wang:2022kia, the analytic expressions produced by the method of this work are much better at capturing the distortions and multimodal features of GW posterior distributions, giving a better balance between accuracy and storage efficiency. The probability-probability (P-P) plot comparing the raw and resampled samples of GW150914 is shown in Fig. 6. The resampling results for other nine GW events are shown in Appendix. B.
To quantify the differences between the distributions of the raw and resampled GW150914 posterior samples, we calculate the Wasserstein distances 1969Markov; computational2019Peyre of the six one-dimensional marginalized distributions between raw and reproduced posteriors. Wasserstein distance is a distance function defined between two probability distributions that measure the similarity of them. It is worth noting that this distance metric characterizes differences in one-dimensional marginal distributions, while corner plots provide a complementary way to compare correlations among parameters. Compared with the more commonly used Kullback-Leibler divergence Kullback:1951zyt, Wasserstein distance is more universal and can reflect the similarity between two distributions which overlap very little. Intuitively, if one regards each probability distribution as an equal (normalized) pile of soil, then the Wasserstein distance can be interpreted as the required minimum “work” to deform one pile of soil into another, which is equal to the amount of soil to be moved multiplied by the average “distance” between two piles (distributions) Wang:2022kia. We normalize the Wasserstein distances using the standard deviation of each one-dimensional marginalized distribution in raw posteriors and the results are shown in Tab. 2. Similar distributions will have a small Wasserstein distance. We can see that most of the one-dimensional marginalized distributions between raw samples and samples resampled with our data products match very well with Wasserstein distances smaller than $0.02$, indicating that the difference is less than $0.02$ standard deviation. Although samples generated with analytic expressions show relatively larger deviation in the marginalized distribution of $\theta_{2}$, analytic expressions outperform the multivariate Gaussian model in all six parameters. As a metric for quantifying differences between high-dimensional distributions, we compute the two-dimensional Hellinger distances hellinger1909neue for pairwise combinations of the parameters of GW150914. The results, summarized in Fig. 7, provide a complementary assessment of the joint correlations.
Additionally, we apply the above KAN-based density estimator to the reconstruction and compression of the full 15-dimensional posterior for GW150914. The data product based on network weights achieves a $34\times$ compression, demonstrating that the method remains robust and effective in the full-parameter scenario. Details of the experimental configuration and the reconstructed posteriors are presented in Appendix C.
Table 1: Comparison between two data-compression products for GW150914’s posterior distribution.GG” denotes the number of Gaussian components in GMMs that fit the conditional probability distribution (1); “Hidden” indicates the number of hidden layers in the neural network architecture before symbolification; “Storage” shows the storage space ratio of data products to raw samples, with the raw data containing 147,634 six-dimensional posterior samples in HDF5 format; “Sampling Time” represents the average re-sampling duration of 20 runs with 2,000 samples per run. Averaged over five runs, network training took 27.14 seconds on the same hardware as in resampling.
Type $G$ Hidden Storage (KB) Sampling Time [^2] NVIDIA RTX 4070 Laptop GPU (8 GB) (s)
Network Weight 5 0 77/6921 0.8809
Analytic Expression 5 0 2.1/6921 0.0900
Table 2: Normalized one-dimensional Wasserstein distances for the posteriors of three data products. Three rows from top to the bottom show the comparisons between raw samples and the samples generated using neural network weights, analytic expressions and fitted multivariate Gaussian distribution, respectively. Each element in the table represents the normalized Wasserstein distance between the one-dimensional marginal distributions.
Type $\mathcal{M}$ $q$ $\chi_{1}$ $\chi_{2}$ $\theta_{1}$ $\theta_{2}$
Network Weight 0.051 0.009 0.008 0.017 0.006 0.019
Analytic Expression 0.050 0.011 0.005 0.006 0.010 0.062
Multivariate Gaussian 0.052 0.261 0.044 0.045 0.134 0.075
Figure 4: Marginalized one- and two-dimensional distributions, comparing samples generated with the neural network weights (orange) and raw samples (blue) of GW150914. Both of them contain 147,634 samples. Contour lines in the two-dimensional joint distributions delineate the $50\%$ and $90\%$ credible regions.
Figure 5: Same as Fig. 4, but for samples generated with analytic expressions (orange), the fitted multivariate Gaussian distribution (grey), and raw samples (blue).
Figure 6: P-P plot comparing the raw posterior samples of GW150914 with those resampled with the trained neural network (left) and fitted analytic expressions (right).
Figure 7: Two-dimensional Hellinger distances for pairwise combinations of the parameters of GW150914. The distances take values in $[0,1]$, with smaller values indicating closer agreement. The top right corner shows the comparison between original samples and samples generated using network weights. The bottom left corner shows the comparison between original samples and samples generated using analytic expressions.
Many key scientific results in GW astronomy, such as constraints on the mass and spin distributions of black holes Dong:2024zzl; Tong:2022iws, rely on population inference that combines posterior samples from many events. In the era of next-generation GW detectors, the number of detected events is expected to increase explosively, and the demand for posterior samples in population studies will grow accordingly. This will pose challenges on users for both data storage and transmission. The primary motivation for proposing a lightweight posterior catalog is to address the need for efficient storage and distribution of large-scale GW posterior data in the future. To ensure that our method preserves not only event-level fidelity but also the statistical properties required for population-level scientific applications, we further validate the reliability of our lightweight data products by performing population inference using samples resampled using our two types of compact data products.
In population inference, we seek to determine the posterior distribution for the population parameters (or hyperparameters), $\bm{H}$, given the collective data $\mathcal{D}=\big\{d_{i}\big\}_{i=1}^{N}$ from $N$ independent events. According to Bayes’ theorem, this posterior, denoted $P(\bm{H}|\mathcal{D})$, is expressed as,
$P(\bm{H}|\mathcal{D})=\frac{P(\mathcal{D}|\bm{H})P(\bm{H})}{P(\mathcal{D})}\,$, (8)
where $P(\mathcal{D}|\bm{H})$ is the total marginalized likelihood over the source parameters of the events and $P(\bm{H})$ is the prior distribution of the population parameters. Using posterior samples in the GW catalog, $P(\mathcal{D}|\bm{H})$ is computed as follows thrane2019introduction,
$P(\mathcal{D}|\bm{H})=\prod_{i=1}^{N}\frac{\mathcal{Z}(d_{i}|\emptyset)}{n_{i}}\sum_{k=1}^{n_{i}}\frac{\pi(\bm{\theta}_{i}^{k}|\bm{H})}{\pi(\bm{\theta}_{i}^{k}|\emptyset)}\,.$ (9)
Here, the summation is over the $n_{i}$ posterior samples, $\{\bm{\theta}_{i}^{k}\}$, for event $i$. The term $\pi(\bm{\theta}_{i}^{k}|\emptyset)$ denotes some default prior in the original analysis and $\mathcal{Z}$ denotes the corresponding evidence, while $\pi(\bm{\theta}_{i}^{k}|\bm{H})$ represents prior in the population model, controlled by the hyperparameters $\bm{H}$.
We compare the posterior distributions, $P(\bm{H}|\mathcal{D})$, based on three sets of samples: the raw samples for each event from GWOSC, the recovered samples using corresponding neural network weights and the recovered samples using corresponding analytic expressions. All three sets contain the same 10 events, and for each event the number of posterior samples equals to that of the raw samples. We use the GWPopulation package Talbot:2024yqw to perform population inference on these three sets of samples. The analyses use the two-component-primary-mass-ratio and iid-spin models, with detailed description and prior choices of the hyperparameters given in Appendix D. Comparative results for the three population-level posterior distributions are shown in Tab. 3 and Fig. 8. As shown in the figure, population inference results of raw samples and samples recovered using neural network weights match very well, validating the feasibility of our compression strategy. Result of samples recovered using analytic expressions exhibits slight deviations. This behavior may in part stem from the numerical instability of hierarchical Bayesian inference Leyde:2023iof, which can amplify small variations in the single-event posteriors, particularly for weakly constrained hyperparameters such as the mass-ratio slope.
Figure 8: Marginalized one- and two-dimensional distributions, comparing the population inference results using raw samples (blue), samples generated with our neural networks (orange) and analytic expressions (pink) of 10 GW events. Contour lines in the two-dimensional joint distributions delineate the $50\%$ and $90\%$ credible regions. The sampling is accomplished using the dynesty sampler Speagle:2019ivv and nlive is set to 2,000.
Table 3: Normalized one-dimensional Wasserstein distances for the population posteriors. The two rows show the comparisons between the population inference results using raw samples and the samples generated using neural network weights and the samples generated using analytic expressions, respectively.
Type $\alpha$ $\beta$ $_{\rm min} mm$$_{\rm max}$ $\lambda_{m}$ $\mu_{m}$ $\sigma_{m}$ $\alpha_{\chi}$ $\beta_{\chi}$ $\xi$ $\sigma$
Network Weight 0.018 0.032 0.027 0.028 0.049 0.012 0.022 0.030 0.015 0.014 0.018
Analytic Expression 0.199 0.225 0.032 0.069 0.169 0.028 0.046 0.165 0.012 0.064 0.215
V Summary
In this work, we propose a KAN-based neural density estimator and explore its potential for constructing high-fidelity, low-storage surrogate GW posterior catalogs. By harnessing the flexibility of KAN’s learnable spline activations, our estimator delivers enhanced interpretability and parameter efficiency. We first demonstrate its analytic-fitting capabilities on verifiable examples and then apply it to real GW posteriors. With a KAN-based estimator, the posterior samples are compactly represented as high-fidelity neural network weights with $\mathcal{O}\,(10^{2})$-fold compression, and analytic expressions with $\mathcal{O}\,(10^{3})$-fold compression, maintaining a balance between fidelity and storage footprint. Once these data products are prepared, users can download these encoded files to flexibly and rapidly resample posterior samples, greatly reducing data-storage and data-transfer overhead. We also validate the fidelity of the compressed data products in a hierarchical population inference, confirming their reliability for downstream high-level analyses.
Our work provides a preliminary demonstration of the parameter efficiency of KAN in density estimation tasks. To the best of our knowledge, this is the first neural density estimator capable of yielding analytic expressions for the PDF. While certain challenges remain in modeling highly complex distributions, the analytic PDF representations obtained by our method have shown clear advantages in storage efficiency and sampling speed. In future work, we plan to further evaluate the performance of the KAN-based neural density estimator in broader application scenarios, such as in GW posterior estimation with conditional networks. This strategy for lightweight GW posterior catalog construction could bring inspiration for tackling the data challenges and pressures of next-generation GW detectors, and it shows how advances in neural network architectures can further empower GW data science.
Acknowledgements
We thank the anonymous referee for comments. This work was supported by the Beijing Natural Science Foundation (1242018), the National Natural Science Foundation of China (123B2043, 12573042), the National SKA Program of China (2020SKA0120300), the Max Planck Partner Group Program funded by the Max Planck Society, and the High-performance Computing Platform of Peking University.
This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan of China.
Code Availability. The code used in this work is publicly available at https://github.com/liu-ws/KMADE.
Appendix A A Case Test Beyond Gaussian Families
In this appendix, we present an additional illustrative example beyond Gaussian families to more comprehensively demonstrate the effectiveness of our method in modeling diverse distributions. In Sec. III, we chose three cases parameterized by Gaussian or Gaussian mixture distributions because they provide closed-form ground truths. It is worth noting that our method remains effective beyond Gaussian families. To illustrate this, we present an additional example using an exponential-type distribution:
$P(\theta_{1},\theta_{2})=\frac{1}{\theta_{1}}e^{-\theta_{1}}e^{-\frac{\theta_{2}}{\theta_{1}}}.$
We fit this target with three mixture components per conditional and obtain the following learned expression,
[…]
$0].\displaystyle+24e^{-0.77(0.65\theta_{1}+0.63\theta_{2}+1.00)^{2}}+00$].
The reconstructed distribution is visualized in Fig. 9. This example demonstrates that, in practice, our estimator can recover the target distribution with high fidelity, serving as an additional validation of the method’s effectiveness.
Figure 9: Results of additional illustrative example. Marginalized one- and two-dimensional distributions are shown in corner plots, comparing samples resampled from fitted expressions (orange) and samples generated from the ground-truth distribution (blue). Both of them contain 40,000 samples.
Appendix B GW Events and Sampling Results
In this appendix, we give some information of binary black hole GW events in GWTC-1 (GW150914, GW151012, GW151226, GW170104, GW170608, GW170729, GW170809, GW170814, GW170818, GW170823 LIGOScientific:2018mvr), and the resampling results obtained with our compact data products.
All ten GW events used in this paper are binary black hole mergers composed of stellar-mass black holes, originally detected during the LIGO/Virgo’s first and second observing runs LIGOScientific:2018mvr. The raw posterior samples used in this paper are taken from the GWTC-2.1 catalog LIGOScientific:2021usb, which features a reanalysis of these events with state-of-the-art waveform models. This population covers a range of total masses from approximately $18.5\,M_{\odot}$ (GW170608) to $84.4\,M_{\odot}$ (GW170729). While most systems have comparable component masses, the spin characteristics show important variations. The majority of events have effective inspiral spins ($\chi_{\rm eff}$) consistent with zero. However, GW151226 stands out with a measured positive spin, providing evidence for at least one spinning black hole in that system. Furthermore, parameter uncertainties vary significantly across the events, often correlating with detection significance: the inaugural loud event GW150914 with a network SNR of $26$ has well-constrained properties, whereas the most distant event GW170729 at redshift $z\approx 0.44$ exhibits much broader posterior distributions due to its lower SNR.
The sampling results are shown in Fig. 10 and Fig.11 and the normalized one-dimensional Wasserstein distances are listed in Tab. 4. As shown in Fig. 10, the samples generated from the neural network weights exhibit close agreement with the original distribution from GWOSC. In Fig. 11, samples generated with the analytic expressions capture the overall structure of the posterior distribution, but exhibit some deviations from the original distributions. In producing the results in the figures, the data preprocessing method, neural network configurations, training processes, and training devices of all ten events are kept consistent. During resampling, the number of samples for each event is the same as that of the raw samples, and the resampling device also remains consistent.
Figure 10: Marginalized one- and two-dimensional distributions, comparing samples generated with the neural network weights (orange) and raw samples (blue) of other nine events. Contour lines in the two-dimensional joint distributions delineate the $50\%$ and $90\%$ credible regions.
Figure 11: Same as Fig. 10, but for samples generated with analytic expressions (orange) and raw samples (blue).
Table 4: Normalized one-dimensional Wasserstein distances for the posteriors of data products of other nine events. Similar to Tab. 2.
Event Type $\mathcal{M}$ $q$ $\chi_{1}$ $\chi_{2}$ $\theta_{1}$ $\theta_{2}$
GW151012 Network Weight 0.012 0.007 0.006 0.008 0.016 0.020
Analytic Expression 0.007 0.011 0.025 0.005 0.016 0.019
GW151226 Network Weight 0.118 0.025 0.010 0.019 0.016 0.036
Analytic Expression 0.116 0.010 0.004 0.004 0.074 0.053
GW170104 Network Weight 0.121 0.014 0.014 0.036 0.021 0.011
Analytic Expression 0.120 0.003 0.005 0.006 0.008 0.069
GW170608 Network Weight 0.066 0.015 0.008 0.014 0.014 0.044
Analytic Expression 0.067 0.018 0.003 0.004 0.020 0.094
GW170729 Network Weight 0.067 0.007 0.007 0.011 0.012 0.009
Analytic Expression 0.022 0.016 0.006 0.007 0.036 0.023
GW170809 Network Weight 0.034 0.018 0.008 0.013 0.012 0.015
Analytic Expression 0.033 0.007 0.008 0.003 0.012 0.051
GW170814 Network Weight 0.031 0.007 0.004 0.015 0.010 0.026
Analytic Expression 0.030 0.005 0.007 0.010 0.008 0.047
GW170818 Network Weight 0.024 0.007 0.009 0.009 0.012 0.024
Analytic Expression 0.022 0.011 0.005 0.006 0.008 0.046
GW170823 Network Weight 0.074 0.012 0.004 0.024 0.023 0.012
Analytic Expression 0.076 0.009 0.005 0.010 0.014 0.010
Appendix C A Full 15-dimensional Posterior Reconstruction for GW150914
In this appendix, we present the reconstruction and compression results for the full 15-dimensional posterior distribution of GW150914 using the KAN-based density estimator. Apart from the input dimension, the network architecture is essentially the same as in the 6-dimensional case, namely a compact two-layer network. Each conditional density is modeled with 5 Gaussian components, and each unmasked edge contains 23 cubic B-spline basis functions. Despite the increased dimensionality, the trained network remains compact. The stored network weights occupy only about 0.5 megabytes, compared with the 16.9 megabytes required to store the original posterior samples, corresponding to a $34\times$ compression ratio. The resulting reconstructed posterior distribution is shown in Fig. 12, which exhibits close agreement with the original distribution, demonstrating the effectiveness and scalability of our KAN-based density estimator.
Figure 12: Marginalized one- and two-dimensional distributions, comparing samples generated with the neural network weights (orange) and raw samples (blue) for the full 15-dimensional posterior of GW150914. Contour lines in the two-dimensional joint distributions delineate the $50\%$ and $90\%$ credible regions. In this figure, we have used $\Delta t_{c}\equiv t_{c}-1126259462$ s.
Appendix D Population Model and Hyperparameters
In this appendix, we provide detailed description and priors of the hyperparameters used in our population inference. We employ the two-component-primary-mass-ratio and iid-spin models from the GWPopulation package Talbot:2024yqw for population inference. The two-component-primary-mass-ratio is a power law model for two-dimensional mass distribution. It models the primary mass and the conditional mass ratio as,
$P(m_{1},q)=$ $P(m_{1})P(q|m_{1})$
$\propto\left[(1-\lambda_{m})m_{1}^{-\alpha}+\lambda_{m}\exp{\left(-\frac{(m_{1}-\mu_{m})^{2}}{2\sigma_{m}^{2}}\right)}\right]\times q^{\beta}\,.$
Here $m_{1}$ is the primary mass, and $q$ is the mass ratio. The iid-spin model assumes independently and identically distributed spins. The magnitudes are assumed to follow a Beta distribution and the orientations are assumed to follow an isotropic and truncated half Gaussian mixture model,
$P(\chi_{1},\chi_{2},\theta_{1},\theta_{1})=$ $P(\theta_{1},\theta_{2})P(\chi_{1},\chi_{2})$
$\propto\left[\frac{(1-\xi)^{2}}{4}+\xi\prod_{i\in{1,2}}\mathcal{N}_{[-1,1]}(\theta_{i};\mu=1,\sigma=\sigma)\right]$
$\times\prod_{j\in[1,2]}{\rm Beta}(\chi_{j};\alpha_{\chi},\beta_{\chi})\,.$
Here $\chi_{1}$ and $\chi_{2}$ are the spin magnitudes of the two black holes, and $\theta_{1}$ and $\theta_{2}$ are the spin orientations. $\rm Beta(\cdot)$ denotes the Beta distribution, $\mathcal{N}_{[-1,1]}$ denotes a truncated Gaussian distribution in the range [1,1][-1,1], and $\alpha_{\chi}$ and $\beta_{\chi}$ are the parameters of the Beta distribution. More description and prior choices of the hyperparameters are shown in Tab. 5. We adopt the prior ranges following the GWPopulation package [^3] https://colmtalbot.github.io/gwpopulation/examples as a test case to validate the effectiveness of our method. More diverse prior choices can be found in previous studies LIGOScientific:2020kqk; KAGRA:2021duu.
Table 5: Definitions and prior distributions for the hyperparameters used in our population inference with the GWpopulation package Talbot:2024yqw. The notation $\mathcal{U}(\theta;a,b)$ represents a uniform prior for the parameter $\theta$ over the interval $[a,b]$.
Notation Parameter description Prior
$\alpha$ Negative power law exponent for the more massive black hole $\mathcal{U}(\alpha;-2,4)$
$\beta$ Power law exponent of the mass ratio distribution $\mathcal{U}(\beta;-4,12)$
$m_{\rm min}$ Minimum black hole mass ($\rm M$$_{\odot}$) $\mathcal{U}(m_{\rm min};2,2.5)$
$m_{\rm max}$ Maximum black hole mass ($\rm M$$_{\odot}$) $\mathcal{U}(m_{\rm max};80,100)$
$\lambda_{m}$ Fraction of black holes in the Gaussian component $\mathcal{U}(\lambda_{m};0,1)$
$\mu_{m}$ Mean of the Gaussian component ($\rm M$$_{\odot}$) $\mathcal{U}(\mu_{m};10,50)$
$\sigma_{m}$ Standard deviation of the Gaussian component ($\rm M$$_{\odot}$) $\mathcal{U}(\sigma_{m};1,10)$
$\alpha_{\chi}$ Parameters of Beta distribution for both black holes $\mathcal{U}(\alpha_{\chi};1,6)$
$\beta_{\chi}$ Parameters of Beta distribution for both black holes $\mathcal{U}(\beta_{\chi};1,6)$
$\xi$ Fraction of black holes in preferentially aligned component $\mathcal{U}(\xi;0,1)$
$\sigma$ Width of preferentially aligned component $\mathcal{U}(\sigma;0.3,4)$